I need help with these questions my dawh1) (10 points) EmKay Industries operates a factory that has fixed costs of $250,000 per month. The variable cost of the produ
Answer #1

As per Cheggs policy 1st question is solved.

Particulars Amount Remarks
Selling Price A             125
Variable Cost B               75
Contribution C=A-B               50
Fixed Cost D    2,50,000
Breakeven Units E=D/C          5,000 Break even Units = Fixed Cost/ Contribution per unit
If Sales are made at 75000 units Remarks
Particulars Amount
Sales F       75,000
Revenue G=F*A 93,75,000
Variable Cost H=F*B 56,25,000
Fixed Cost D    2,50,000
Total Cost I=D+H 58,75,000
Net Profit K=G-I 35,00,000
Average Cost Per Unit J=I/F         78.33 Average cost per unit = Total cost/ number of units
Marginal Cost P.U               75 Equals to Variable cost per unit

Please Upvote and Support!!.

Thanks